AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 303     Entry time: Sun Feb 27 12:00:13 2022
Author: RSS TD 
Subject: Current status 

Ring current ~7.7e7 particles/s

Merger, Tape server (WS3): 1414 kB/s writing ok

Leakage current (WS4): 3.43 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

Attachment 1: 13-xy-2022-02-27_13-01-36.png  269 kB  | Hide | Hide all
13-xy-2022-02-27_13-01-36.png
Attachment 2: 13-yrate-2022-02-27_13-01-06.png  156 kB  | Hide | Hide all
13-yrate-2022-02-27_13-01-06.png
Attachment 3: 13-xrate-Screenshot_from_2022-02-27_13-00-50.png  156 kB  | Hide | Hide all
13-xrate-Screenshot_from_2022-02-27_13-00-50.png
ELOG V3.1.4-unknown